1) When getting into shape by exercising, the subject’s maximum recommended number
of heartbeats per minute (h) can be determined by subtracting the subject’s age (a)
from 220 and then taking 75% of that value. This relation is expressed by which of the
following formulas?
A. h = .75(220 – a)
B. h = .75(220) – a
C. h = 220 – .75a
D. .75h = 220 – a
E. 220 = .75(h – a)

2)An airplane flew for 8 hours at an airspeed of x miles per hour (mph), and for 7 more
hours at 325 mph. If the average airspeed for the entire flight was 350 mph, which of
the following equations could be used to find x ?
A. x + 325 = 2(350)
B. x + 7(325) = 15(350)
C. 8x – 7(325) = 350
D. 8x + 7(325) = 2(350)
E 8x + 7(325) = 15(350)

Answer by mananth(16946)

1) When getting into shape by exercising, the subject’s maximum recommended number
of heartbeats per minute (h) can be determined by subtracting the subject’s age (a)
from 220 and then taking 75% of that value. This relation is expressed by which of the
following formulas?
A. h = .75(220 – a)
B. h = .75(220) – a
C. h = 220 – .75a
D. .75h = 220 – a
E. 220 = .75(h – a)
..
A. h = 0.75(220 – a)
2)An airplane flew for 8 hours at an airspeed of x miles per hour (mph), and for 7 more
hours at 325 mph. If the average airspeed for the entire flight was 350 mph, which of
the following equations could be used to find x ?
.
8x+2275 /15 = 350
8x+7*325 = 350*15
Answer = E
A. x + 325 = 2(350)
B. x + 7(325) = 15(350)
C. 8x – 7(325) = 350
D. 8x + 7(325) = 2(350)
E 8x + 7(325) = 15(350)